A town has a population of 2000 and grows at 4% every year. What will be the population after 15 years, to the nearest whole number?



1
Expert's answer
2022-04-15T09:09:50-0400

"x\left( n \right) -population\,\,after\,\,n\,\,years\\x\left( n \right) =x\left( 0 \right) \cdot 1.04^n=2000\cdot 1.04^n\\x\left( 15 \right) =2000\cdot 1.04^{15}=3601.89\approx 3602"